40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 11 00:07:54 2015, rana, Update, LSC, Lock Loss plot 1107673198.png1107673198.png
    Reply  Wed Feb 11 02:42:05 2015, Jenne, Update, LSC, PRC error signal RF spectra REFL33_25.pdfREFL33_15.pdf
       Reply  Wed Feb 11 03:41:12 2015, Koji, Update, LSC, PRC error signal RF spectra 
          Reply  Wed Feb 11 17:31:11 2015, ericq, Update, LSC, RFPD spectra REFL.pngAS.pngPOP.png2015-02-PDSpectra.zip
             Reply  Wed Feb 11 18:07:42 2015, ericq, Update, LSC, RFPD spectra 
                Reply  Thu Feb 12 01:00:18 2015, rana, Update, LSC, RFPD spectra 
Message ID: 10998     Entry time: Wed Feb 11 00:07:54 2015     Reply to this: 10999
Author: rana 
Type: Update 
Category: LSC 
Subject: Lock Loss plot 

Here is a lock loss from around 11 PM tonight. Might be due to poor PRC signals.  \oint {\frac{\partial PRCL}{\partial x}}

This is with arm powers of ~6-10. You can see that with such a large MICH offset, POP22 signal has gone done to zero. Perhaps this is why the optical gain for PRCL has also dropped by a factor of 30 crying.

This seems untenable no. We must try this whole thing with less MICH offset so that we can have a reasonable PRCL signal.cool

Attachment 1: 1107673198.png  460 kB  | Hide | Hide all
1107673198.png
ELOG V3.1.3-